TMS Meeting Minutes, April 15, 2021

Attendees: Andrew, Breann, Heejoo, John, Olga, Trenton, Yang (Zoom Meeting)

Heejoo's Mountain Visit
  • WELL DONE, HEEJOO!
  • Was able to compress a planned three day trip into two days (no conflicting scheduling on the telescope)
  • Achieved nice alignment of collimators and retros on SX as evidenced by Etalon channel alignment GUI
  • By increasing the gain, can see 9 channels (standard gain gives 8 channels)
    • Andrew recommends setting the gain so that the signal strength is at 30% at Zenith (providing for gain residual)
  • At 63.5 deg elevation, Ch. 4 and Ch. 22 have low fiber output (likely a fiber issue), while Ch. 11 and Ch. 18 appear to have an unstable mount (low power at 63.5 deg)
    • By simply touching the collimator, the channel return increases substantially
    • It appears there is a a bimodal positioning issue with the collimator alignment for these channels
  • Completed laser tracker survey on the SX side, though the data has yet to be processed
    • Heejoo still has to extend the vectors from the survey to find the fiber tip positions relative to the measured collimator positions
    • Andrew and Heejoo will be working on computing the positional datasets this week
    • Andrew suggests setting the origin by calculating the conicoid vertex and setting the offset of the retro surface
    • If we know the primary mirror prescription, the origin can be set to the conicoid vertex in the center of the M1 hole
    • Must take into account the offset of the SMR vertex off of the surface of the mirror by a certain amount in Z, dropped down as necessary from the edge of the mirror
    • Don't yet have the LBTI and primary diameter retros measured at this time
  • Solid retro test (42 meter propagation length)
    • Thorlabs
      • Inc. Ang. 0 -> 10% return
      • Inc. Ang. 20 -> 7% return
      • Inc. Ang. 30 -> 0% return (could be a spot size issue)
    • Wide-angle
      • Inc. Ang. 0 -> 9% return
      • Inc. Ang. 20 -> 8% return
      • Inc. Ang. 30 -> 0% return (could be a spot size issue)
    • Shows that this retro works at least as well as a standard retro, works much better for wide-angles and closer distances
Bino-TMS Code
  • Not quite ready yet
  • Yang thinks there should be no problem interfacing with John's wrapper program
  • The main changes will be in the measurement and pose calculation code
  • Yang needs some time to make the necessary changes
  • Requests the new laser tracker data set to be used in preparing the code
  • Andrew thinks we will be ready for bino code next week
  • Once the laser tracker post processing is completed, we will be ready
Thorlabs Collimator
  • Thorlabs had a sign error in their fabrication
  • The wrong optical spec hardware should arrive tomorrow
  • Will take two more weeks to get the corrected part
  • Everyone makes mistakes, it seems
Multiline Software Reconstruction Proposal
  • Specific TCP/IP controls called out, status updates, improved overall functionality, etc.
  • Still waiting for response from Etalon software team about the crashing problems
  • Etalon did find some correlation to one of the port errors
Thermal Telemetry Correlation to M1 Focus
  • Front to back thermal gradient has the most direct effect on thermal focus
  • Can we take the bulk of measured Z4 drift and fit the thermal fluctuations on mirror
  • Compare thermocouple data from 4/8/2021
  • Could be interesting to run the reduction again and calculate
  • Very first time we have been able to remove structural thermal focus changes and look just
  • Once we have these results, can go and check this against all of our passive data
  • This thermal model can be used to update the glass thermal focus fit for the LUT collimation
  • Reliable thermal focus for the primary mirror
  • TMS maintaing position, thermal model correcting glass changes
  • Basic noise in FPIA and instability produces a wavelength floor for LBTI
  • LBTI can operate somewhat reliably at 1 um, but improvements no TMS and FPIA could push this into the VIS
LBC Scheduling
  • During the next four nights, there are plans to use LBC's at some point, though we don't know for sure when
  • Big dataset two nights ago (pull active optics Zernikes)

-- TrentonBrendel - 15 Apr 2021
Topic revision: r1 - 15 Apr 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback